# Hypergraph Neural Network with State Space Models for Node Classification

### v1

@@ Abstract
In recent years, graph neural networks (GNNs) have gained significant attention for node classification tasks on graph-structured data. However, traditional GNNs primarily focus on adjacency relationships between nodes, often overlooking the rich role-based characteristics that are crucial for learning more expressive node representations. Existing methods for capturing role-based features are largely unsupervised and fail to achieve optimal performance in downstream tasks. To address these limitations, we propose a novel hypergraph neural network with state space model (HGMN), that effectively integrates role-aware representations into GNNs and state space model. HGMN utilizes hypergraph construction techniques to model higher-order relationships and combines role-based and adjacency-based representations through a learnable mamba transformer mechanism. By leveraging two distinct hypergraph construction methods—based on node degree and neighborhood levels—it strengthens the connections among nodes with similar roles, enhancing the model’s representational power. Additionally, the inclusion of hypergraph convolution layers enables the model to capture complex dependencies within hypergraph structures. To mitigate the over-smoothing problem inherent in deep GNNs, we incorporate a residual network, ensuring improved stability and better feature propagation across layers. Extensive experiments conducted on one newly introduced dataset and four benchmark datasets demonstrate the superiority of HGMN. The model achieves significant performance improvements on node classification tasks compared to state-of-the-art GNN methods. These results highlight HGMN’s ability to provide enriched node representations by effectively embedding role-based features alongside adjacency information, making it a versatile and powerful tool for a variety of graph-based learning applications.
@@ 1 Introduction
Graph-based modeling has become a prevalent approach for analyzing complex data structures and relationships, including applications like social networks [1], molecular interactions [2], and brain connectivity [3]. Graph neural networks (GNNs) have demonstrated superior performance over traditional neural networks in a variety of applications [4]. Prominent examples of GNNs include models such as GCN [5], GAT [6], and GraphSAGE [7]. These networks are designed to capture the spatial characteristics inherent in graph data, enhancing both node-level and graph-level representations. A key principle underlying most GNNs is the aggregation of information from neighboring nodes, a concept often referred to as message passing [8]. This approach focuses on leveraging features from adjacent or connected nodes within the graph to improve representation learning.
In a graph, nodes typically represent different social roles or functions, and nodes that share similar roles or functions are not always directly connected. Traditional GNNs, such as GCN [5], GAT [6], and GraphSAGE [7], generate feature embeddings for central nodes by aggregating information from their neighbors. While this approach captures local subgraph structures, it has notable limitations: (1) It faces challenges in generating node representations that capture detailed and role-specific features effectively, and (2) It encounters difficulties in linking nodes with similar roles that are distant or lack direct connections. A node’s role can reflect its real-world semantic function such as being a student, advisor, or manager, or its structural role within a network, such as an opinion leader or structural hole spanner [9]. Although unsupervised methods have been proposed for role representation learning [10, 11], their performance in downstream tasks remains suboptimal [7]. Furthermore, adjacency-based multi-layer GNNs frequently face over-smoothing problems, where nodes become indistinguishable after several layers of message passing.
Traditional GNNs primarily work with standard graphs that represent node adjacency while overlooking the role-based relationships between nodes. In contrast, hypergraphs [12] can be derived from conventional graph structures in multiple ways. Hypergraph adjacency encodes richer relationships among nodes and clusters, enhancing the effectiveness of aggregation processes in graph neural networks (NNs) for capturing both role-based and adjacency-based representations.
Recently, Graph Transformers have become a powerful tool due to their ability to effectively capture long-range interactions between nodes [13, 14]. Unlike traditional local message-passing methods that focus on edge-level neighborhood structures [5, 15], the attention mechanisms in Graph Transformers allow each node to interact with every other node in the graph, providing a more global perspective [16]. To enhance the development of Graph Transformers, the GraphGPS framework integrates structural/positional encodings (SE/PE), a message-passing NN (MPNN), and an attention module. Collectively, these components refine node and edge embeddings to facilitate downstream tasks. This modular design allows for the seamless incorporation of various attention mechanisms, offering significant flexibility for users [17].
@@
The quadratic computing cost of the attention mechanism limits the usefulness of transformers for handling long sequences, despite their significantly enhanced computational capability. Furthermore, it might not always be optimal for long-range dependencies to fully capture attention or encode all contextual information. Empirical studies suggest that increasing context length often fails to improve the performance of many sequence models [18]. State space models (SSMs) have surfaced as an effective alternative to attention-based frameworks such as Transformers [19]. However, their performance falls short of Transformers due to constraints in capturing context compression that depends on the input, a drawback of their time-invariant transition mechanisms. Mamba has shown remarkable success in language modeling, exceeding the performance of Transformers of the same size and achieving results comparable to Transformers that are twice as large. These achievements have inspired recent efforts to adapt Mamba’s architecture for various data modalities [20, 21].
Building on the concept of hypergraphs [12], we introduce the hypergraph NN with state space model (HGMN), which integrates role-based representations into GNNs. HGMN leverages hypergraph construction methods to integrate node representations of roles and adjacency from higher-order viewpoints. We incorporate two methods for constructing hypergraphs that aggregate node information based on degree and neighborhood levels, reinforcing connections between nodes with similar roles within the hypergraph. HGMN integrates a mamba transformer mechanism with adaptive weights to merge adjacency-based and role-based representations. Through hypergraph convolution layers, the model learns hypergraph structures during training. A residual network is used to address the over-smoothing problem between convolutional layers. Compared to standard GNNs, HGMN enriches node representations by incorporating role characteristics and adopts a more effective aggregation approach. It is versatile and well-suited for node representation tasks involving both adjacency structures and role information.
@@
    We propose a novel hypergraph NN with state space model (HGMN), that integrates role-based representations into graph NNs. HGMN leverages hypergraph construction techniques to effectively combine adjacency and role-based node representations from higher-order perspectives.
@@
    By employing hypergraph convolution layers, HGMN learns hypergraph structures during training. To address the over-smoothing problem common in graph NNs, a residual network is incorporated, improving the stability and expressiveness of the model.
@@
    HGMN effectiveness is validated through comprehensive experiments on four publicly available datasets. The results demonstrate that HGMN achieves significant performance gains, with absolute improvements of up to $12.1\%$ in node classification accuracy compared to existing GNN methods.
@@ 2 Preliminaries and related work
An undirected graph can be represented as $G=(V,E,A)$, where $V=\{v_{1},v_{2},\dots,v_{N}\}$ is the set of nodes, $A\in\mathbb{R}^{N\times N}$ is the adjacency matrix, and $E$ represents the set of edges. The role representation is denoted as $X_{r}\in\mathbb{R}^{N\times F_{r}}$, while $X_{a}\in\mathbb{R}^{N\times F_{a}}$ corresponds to the adjacency-based representation. Node labels are indicated by $Y\in\mathbb{R}^{N}$, where the nodes belong to $M$ distinct categories. Graph neural networks (GNNs) generate the node representations $Z\in\mathbb{R}^{N\times F_{h}}$ and use them to predict the labels of the nodes.
@@ 2.1 Graph Neural Network
Graph Neural Networks (GNNs), introduced in 2009, are designed to analyze data that is structured as graphs [22]. Over recent years, research interest in GNNs has grown significantly [23, 24]. This research can generally be divided into two main categories. The first focuses on applying GNNs to a variety of tasks, including knowledge representation [25], text generation [26], traffic prediction [27], and weather forecasting [28].
ChebNet [29] introduced a spectral approach using Chebyshev polynomials to extract features from graph-structured data. Building on this, GCN [5] simplified the method with feature propagation to aggregate adjacency information, improving computational efficiency. Shi et al. [30] introduced UniMP, a consolidated message-passing framework that combines feature and label propagation during both the training and inference stages. SAT [14] enhances the performance of existing GNNs by extracting subgraph representations, improving their overall effectiveness. These GNNs primarily rely on adjacency-based message-passing for feature aggregation.
Beyond adjacency-driven feature aggregation, some GNNs incorporate additional characteristics such as node identities or positional information, which are absent in traditional methods. For instance, ID-GNN [31] adopts an inductive approach to incorporate node identities into the message-passing process. Similarly, the position-aware GNN (PGNN) [32] focuses on generating embeddings that account for node locations. Li et al. [33] introduced GLOGNN and its enhanced version, GLOGNN++, which utilizes a coefficient matrix to enable neighborhood aggregation and capture inter-node correlations.
Xie et al. [34] developed the scale-aware GNN (SAGNN), which models cross-scale interactions for tasks such as few-shot semantic segmentation. At the same time, Xu et al. [35] introduced the graph wavelet NN (GWNN), which utilizes graph wavelet transforms to overcome the limitations of earlier spectral graph convolutional networks that depend on graph Fourier transforms. Although these approaches demonstrate strong performance, they do not directly tackle the connection between role-based information and higher-order feature aggregation within graph structures.
@@ 2.2 Hypergraph Learning
Hypergraph learning was first proposed as a propagation method on hypergraph structures. Thanks to its degree-free hyperedges, hypergraph learning enables the modeling of high-order data correlations beyond simple pairwise relationships [36]. Early works, such as those by Feng et al. [12] and Yadati et al. [37], were pioneers in integrating hypergraph structures into GNNs, marking a significant step in the evolution of GNNs. Bai et al. [38] advanced this direction by introducing hypergraph attention networks and hypergraph convolutional NNs to enhance representation learning capabilities. Zhang et al. [39] proposed the dual-channel hypergraph convolutional framework (DHCF) for collaborative filtering in recommendation systems, demonstrating hypergraphs’ utility in practical applications. Chien et al. [40] contributed a generalized framework for hypergraph NNs that dynamically chooses propagation methods suited to the characteristics of individual datasets. Yang et al. [41] further refined this concept by introducing an attention-based hypergraph NN that prioritizes essential features during learning. Despite these advancements and the broad application of hypergraph NNs across various domains, their potential for handling node representations rich in role-based features remains largely unexplored.
@@ 2.3 State Space Models
State space models (SSMs), which range from hidden Markov models to recurrent NNs (RNNs) in contemporary deep learning, use recurrent updates over sequences through hidden states. Using a recurrent scanning approach, SSMs maintain context within hidden states and update outputs by integrating these states with incoming inputs. structured state space models (S4) improve computational efficiency by leveraging reparameterization techniques, providing a streamlined alternative to traditional attention mechanisms [42].
Recent developments in linear-time attention models based on S4 include innovations such as hungry hungry hippos (H3) [43], hyena [44], and receptance weighted key value RWKV [45]. Building on this foundation, Mamba integrates a data-driven selection mechanism within S4, allowing for efficient capture of long-range dependencies as the sequence lengths increase [18]. Mamba achieves linear-time complexity in long-sequence modeling and outperforms Transformers across various benchmarks. Moreover, its adaptability extends beyond sequential data, demonstrating effectiveness in tasks like image segmentation by enhancing long-range contextual understanding [46].
Figure 1: Overview of the proposed framework
Figure 2: Simple graph vs. Hypergraph. (a) A simple graph along with its vertex-edge incidence matrix. (b) A hypergraph and its corresponding vertex-edge incidence matrix.
@@ 3 Hypergraph Neural Network with State Space Model (HGMN)
Figure 1 depicts the structure of the HGMN framework, comprising three primary components. Initially, the graph structure is encoded to derive the role and adjacency representations using Node2vec [12] and Graphwave [11], respectively. A hypergraph is then constructed from the input graph. These representations are combined into a unified embedding through an attention mechanism. Finally, a multi-layer hypergraph convolutional network utilizes the fused embeddings and hypergraph adjacency matrix for iterative computations.
@@ 3.1 Construction of Hypergraph
This subsection outlines two detailed methods for constructing hypergraphs. The resulting hypergraphs are utilized in the following hypergraph convolutional networks to perform feature aggregation.
3.1.1 Description of the Hypergraph
In a standard graph, edges connect pairs of nodes, while in a hypergraph, a hyperedge can connect several nodes that share common characteristics. This allows a simple graph to be restructured into a hypergraph. A hypergraph $G_{H}=(V_{H},E_{H},W_{H})$ is defined with $H$ represented as $\theta(v,e)$, where $\theta(v,e)$ is either 0 or $1$, signifying whether node $v$ belongs to a hyperedge $e$. $V_{H}$ denotes the set of nodes, $E_{H}$ represents the set of hyperedges, and $W_{H}$ is the hyperedge weight matrix. The degree of a hyperedge is given by:
$d(e_{i})=\sum_{v_{i}\in V}\theta(v_{i},e_{i}).$ (1)
The node degree is computed as:
@@ 3.1.1 Description of the Hypergraph
where the degree matrices for nodes and hyperedges are given by $D_{v}$ and $D_{e}$, respectively. The techniques for hypergraph construction in HGMN are shown in Fig.2.

### v2

@@ Abstract
In recent years, graph neural networks (GNNs) have gained significant attention for node classification tasks on graph-structured data. However, traditional GNNs primarily focus on adjacency relationships between nodes, often overlooking the role-based characteristics that can provide complementary insights for learning expressive node representations. Existing frameworks for extracting role-based features are largely unsupervised and often fail to translate effectively into downstream predictive tasks. To address these limitations, we propose a hypergraph neural network with a state space model (HGMN). The model integrates role-aware representations into GNNs by combining hypergraph construction with state-space modeling in a principled manner. HGMN employs hypergraph construction techniques to capture higher-order relationships and leverages a learnable mamba transformer mechanism to fuse role-based and adjacency-based embeddings. By exploring two distinct hypergraph construction strategies, degree-based and neighborhood-based, the framework reinforces connectivity among nodes with structural similarity, thereby enriching the learned representations. Furthermore, the inclusion of hypergraph convolution layers enables the model to account for complex dependencies within hypergraph structures. To alleviate the over-smoothing problem encountered in deeper networks, we incorporate residual connections, which improve stability and promote effective feature propagation across layers. Comprehensive experiments on benchmark datasets including OGB, ACM, DBLP, IIP TerroristRel, Cora, Citeseer, and Pubmed demonstrate that HGMN consistently outperforms strong baselines in node classification tasks. These results support the claim that explicitly incorporating role-based features within a hypergraph framework offers tangible benefits for node classification tasks.
@@ 1 Introduction
Graph-based modeling has become a powerful paradigm for analyzing complex relational data, with applications spanning social networks [1], molecular interactions [2], and brain connectivity [3]. Graph neural networks (GNNs) extend traditional neural architectures by leveraging graph structure, achieving strong performance across diverse tasks [4]. Representative models such as GCN [5], GAT [6], and GraphSAGE [7] exemplify this progress. These frameworks operate under the message-passing framework [8], where node representations are iteratively updated by aggregating features from their neighbors. This design effectively captures local structural dependencies and enhances both node- and graph-level representations. In real-world graphs, nodes often represent distinct roles or functions, yet nodes with similar roles are not always directly connected. Traditional GNNs such as GCN [5], GAT [6], and GraphSAGE [7] generate embeddings by aggregating information from immediate neighbors. Jumping knowledge networks (JK-Nets) [9] aggregate information across GNN layers to adaptively capture neighborhood features at different depths. While effective for capturing local structures, this framework has two major limitations: (1) it struggles to produce role-specific representations that capture fine-grained semantic or structural functions, and (2) it fails to connect role-similar nodes that are distant or lack direct edges. A node’s role may reflect semantic attributes (e.g., student, advisor, manager) or structural properties within the network (e.g., opinion leader, bridge, structural hole spanner) [10]. Although unsupervised role representation frameworks [11, 12] have been proposed, their effectiveness in downstream tasks remains limited. Moreover, adjacency-based multi-layer GNNs are prone to over-smoothing, where node embeddings become indistinguishable after repeated message passing. Traditional GNNs mainly operate on standard graphs that capture pairwise node adjacencies, often neglecting the role-based relationships that exist among nodes. In contrast, hypergraphs [13] can be constructed from conventional graphs in various ways, enabling the modeling of higher-order interactions. By encoding richer connections among nodes and clusters, hypergraph adjacency facilitates more expressive aggregation, allowing graph neural networks to effectively integrate both role-based and adjacency-driven representations.
Recently, graph transformers have emerged as powerful models for capturing long-range interactions between nodes [14, 15]. Unlike traditional message-passing frameworks that emphasize local edge-level neighborhoods [5, 16], attention mechanisms in graph transformers enable each node to attend to all others, offering a global perspective on graph structure [17]. To further advance this paradigm, the GraphGPS framework combines structural and positional encodings (SE/PE), a message-passing neural network (MPNN), and an attention module. Together, these components enrich node and edge embeddings and improve performance on downstream tasks. Its modular design also supports the seamless integration of diverse attention mechanisms, providing flexibility and adaptability for different applications [18]. Graphormer [19] adapts transformers to graphs, capturing long-range dependencies, but standard transformers often underperform on graph-specific tasks due to limited structural awareness. GROVER [20] addresses this with self-supervised pretraining on large-scale molecular graphs, learning richer node and motif representations. GraphMAE [21] leverages masked graph autoencoding to capture both local and global patterns efficiently.
@@
The quadratic computational cost of the attention mechanism restricts the scalability of transformers for long sequences, despite their strong modeling capacity. Moreover, fully attending to all tokens is not always necessary or optimal, as long-range dependencies may not require exhaustive contextual encoding. Empirical evidence indicates that simply increasing context length often fails to yield consistent improvements in sequence model performance [22]. State space models (SSMs) have emerged as a promising alternative to attention-based architectures [23], offering greater efficiency for long sequences. However, their effectiveness is limited by time-invariant transition mechanisms, which constrain their ability to perform input-dependent context compression. Recently, Mamba has demonstrated remarkable success in language modeling, surpassing transformers of the same size and achieving performance on par with models that are twice as large. These advancements have motivated growing interest in extending Mamba’s architecture to diverse data modalities [24, 25].
Building on the concept of hypergraphs [13], we propose the hypergraph neural network with a state space model (HGMN). This framework is designed to integrate both adjacency-based and role-based representations into graph learning. Unlike conventional GNNs that primarily rely on local message passing, HGMN employs two complementary hypergraph construction strategies based on node degree and neighborhood levels. These strategies capture higher-order relationships and reinforce connections among nodes with similar roles, even when they are distant in the original graph. To effectively unify these heterogeneous perspectives, HGMN incorporates a Mamba-inspired state space mechanism with adaptive weighting. The mechanism enables a flexible balance between structural and role-driven features. Hypergraph convolution layers are further utilized to jointly learn adjacency and role-based dependencies, while residual connections mitigate over-smoothing and enhance stability during training. Collectively, these design choices make HGMN a robust and versatile model, capable of producing richer and more discriminative node representations than standard GNNs. As a result, graph learning tasks that require the integration of both structural connectivity and role-aware semantics are significantly enhanced.
@@
    We propose a novel hypergraph neural network with state space model (HGMN), which integrates role-based representations into graph neural networks. HGMN leverages hypergraph construction techniques to effectively combine adjacency and role-based node representations from higher-order perspectives.
@@
    By employing hypergraph convolution layers, HGMN learns hypergraph structures during training. To address the over-smoothing problem common in graph neural networks, a residual network is incorporated, improving the stability and expressiveness of the model.
@@
    HGMN effectiveness is validated through comprehensive experiments on four publicly available datasets. The results demonstrate that HGMN achieves significant performance gains, with absolute improvements of up to $12.1\%$ in node classification accuracy compared to existing GNN frameworks.
Table 1: List of symbols and their definitions used in HGMN.
Symbols Definition and description
$G$ A general undirected graph
$E_{l},E_{d}$ Hyperedge by node link, and node degree
$X_{r},X_{a}$ Role embeddings, and adjacency embeddings
$G_{h}^{l},G_{h}^{d}$ Hypergraphs constructed by node link, and node degree
$D_{v}^{l},D_{v}^{d}$ Node degree matrices by node link, and node degree
$D_{e}^{l},D_{e}^{d}$ Hyperedge degree matrices by node link, and node degree
$Y$ The label of nodes
$J$ The embedding of nodes
@@ 2 Preliminaries and related work
An undirected graph can be represented as $G=(V,E,A)$, where $V=\{v_{1},v_{2},\dots,v_{N}\}$ is the set of nodes, $A\in\mathbb{R}^{N\times N}$ is the adjacency matrix, and $E$ represents the set of edges. The role representation is denoted as $X_{r}\in\mathbb{R}^{N\times F_{r}}$, $X_{a}\in\mathbb{R}^{N\times F_{a}}$ corresponds to the adjacency-based representation, and $X_{f}\in\mathbb{R}^{N\times F_{h}}$ denotes the fused embedding. Node labels are indicated by $Y\in\mathbb{R}^{N}$, where the nodes belong to $M$ distinct categories. GNNs generate the node representations $J\in\mathbb{R}^{N\times F_{h}}$ and use them to predict the labels of the nodes. For hypergraph construction, the incidence matrix is denoted as $H$, with $D_{v}$ and $D_{e}$ representing the node degree and hyperedge degree matrices, respectively. Model training involves learnable parameters $\delta$, and the coefficients $\lambda$, $\alpha$, and $\beta$ are used for regularization and feature fusion. Table 1 lists the notations used in the HGMN model.
@@ 2.1 Graph Neural Network
Graph Neural Networks (GNNs), first introduced by Scarselli et al. [26], are designed to analyze graph-structured data and have achieved success in diverse domains such as knowledge representation [27], text generation [28], traffic prediction [29], and weather forecasting [30]. Early spectral frameworks, including ChebNet [31] and its simplification GCN [5], pioneered localized feature propagation. Later frameworks, such as UniMP [32] and SAT [15], further extended adjacency-based message passing by incorporating label propagation or subgraph representations.
To enrich graph representations beyond adjacency, models such as ID-GNN [33] and PGNN [34] integrated node identity and positional encodings, while GLOGNN and GLOGNN++ [35] introduced coefficient matrices to capture higher-order correlations. Other directions explored multiscale and spectral perspectives, for example, SAGNN [36] models cross-scale interactions, and GWNN [37] leverages wavelet transforms to address limitations of Fourier-based methods. Recent progress in graph representation learning has demonstrated broad applicability across domains such as fraud detection, multi-agent dynamics, fault diagnosis, and intelligent resource management [38]. The H2IDE model [39] addresses homophily–heterophily challenges in multi-relational graphs using disentangled representation learning and relation-aware attention for effective fraud detection. Extending structural modeling, Shi et al. [40] employed hypergraphs and a Fokker–Planck framework to capture higher-order interactions in multi-agent systems. Interpretability has also gained prominence, with Wang et al. [41] introducing a causal graph, a guided GNN for interpretable fault diagnosis in nuclear systems. Similarly, the dynamic parking model [42] leverages neural prediction and control for adaptive resource allocation.
@@ 2.2 Hypergraph Learning
Hypergraph learning was first proposed as a framework for propagating information over hypergraph structures, where degree-free hyperedges enable modeling of high-order correlations beyond simple pairwise relationships [43]. Building on this foundation, Feng et al. [13] and Yadati et al. [44] pioneered the integration of hypergraph structures into GNNs, laying the groundwork for subsequent advancements. For example, Bai et al. [45] introduced hypergraph attention networks and hypergraph convolutional neural networks to strengthen representation learning, while Zhang et al. [46] developed the dual-channel hypergraph convolutional framework (DHCF) to improve collaborative filtering in recommendation systems. Further generalizations have been proposed, such as Chien et al. [47], who introduced a framework that dynamically selects propagation methods according to dataset characteristics, and Yang et al. [48], who designed an attention-based hypergraph neural network to prioritize informative features. Collectively, these works establish hypergraph neural networks as powerful tools for capturing complex, higher-order dependencies across domains. However, their potential for explicitly modeling role-based node representations, which are crucial for tasks where adjacency alone is insufficient, remains largely unexplored.
@@ 2.3 State Space Models
State space models (SSMs) form a broad family of sequence modeling frameworks, ranging from classical hidden Markov models to recurrent neural networks (RNNs) in modern deep learning. These models maintain context by recurrently updating hidden states, integrating new inputs with past information to generate outputs. While effective for sequential data, traditional SSMs often face challenges in scalability and efficiency. To address these limitations, structured state space models (S4) introduced reparameterization techniques that significantly improve computational efficiency, offering a streamlined alternative to the quadratic cost of attention mechanisms [49].
Building on S4, a wave of linear-time attention models has been proposed, including hungry hungry hippos (H3) [50], Hyena [51], and receptance weighted key value (RWKV) [52], all designed to capture long-range dependencies while maintaining efficiency. Extending this line of work, Mamba introduces a data-driven selection mechanism within S4, enabling adaptive context compression as sequence lengths grow [22]. This design achieves linear-time complexity while outperforming Transformers of comparable or larger sizes across diverse long-sequence benchmarks. Moreover, Mamba’s adaptability has been demonstrated beyond sequential domains, including computer vision tasks such as image segmentation, where its ability to capture long-range contextual information leads to substantial performance gains [53].
Figure 1: Overview of the proposed HGMN. The input graph is first converted into a hypergraph to capture higher-order node relationships. Node features are transformed into adjacency-based ($X_{a}$) and role-based ($X_{r}$) embeddings, which are fused via the Mamba module using linear projections, convolution, and a state-space model, producing $X_{f}$. Fused embeddings pass through hierarchical hypergraph convolution layers with downsampling to generate multi-level representations ($X_{1},\dots,X_{last}$), which are aggregated and fed to a classifier for node-level prediction.
Figure 2: Simple Graph vs. Hypergraph. (a) A traditional graph with edges connecting pairs of nodes and its vertex-edge incidence matrix. (b) A hypergraph where hyperedges can connect multiple nodes simultaneously, with the corresponding incidence matrix. Hypergraphs capture higher-order relationships that simple graphs cannot, motivating the proposed HGMN to model complex node interactions for improved representation learning.
@@ 3 Hypergraph Neural Network with State Space Model (HGMN)
The proposed HGMN framework captures both role-based and adjacency-based node features while modeling long-range dependencies by integrating hypergraph representation learning with state-space models. To illustrate the overall workflow, Fig. 1 outlines four main stages: first, role and adjacency features are extracted from the input graph using Node2vec [13] and Graphwave [12]; next, hypergraphs are constructed to encode higher-order relationships among nodes; these features are then adaptively fused using a Mamba-based structured state-space model (SSM) attention mechanism; finally, the fused representations are propagated through hypergraph convolutional layers, enhanced via residual learning, and fed into the classification module. This staged design ensures that HGMN effectively integrates local, global, and higher-order structural information for robust graph representation learning.
@@ 3.1 Construction of Hypergraph
To extend traditional pairwise graphs into higher-order structures, HGMN relies on hypergraph construction. Unlike a standard graph, where an edge connects only two nodes, a hypergraph allows a hyperedge to connect multiple nodes that share common characteristics, thereby capturing richer relationships among them. Formally, a hypergraph is defined as $G_{H}=(V_{H},E_{H},W_{H})$, where $V_{H}$ denotes the set of nodes, $E_{H}$ the set of hyperedges, and $W_{H}$ the hyperedge weight matrix. The incidence relation is described by $\theta(v,e)\in\{0,1\}$, which indicates whether node $v$ belongs to hyperedge $e$. Based on this definition, the degree of a hyperedge and node is given by:
$d(e_{i})=\sum_{v_{i}\in V}\theta(v_{i},e_{i}),$ (1)
@@
with the corresponding degree matrices for nodes and hyperedges denoted by $D_{v}$ and $D_{e}$. These formulations establish the mathematical foundation for constructing hypergraphs within HGMN, as illustrated in Fig. 2.
To capture different structural perspectives, we employ two complementary strategies for constructing hypergraphs: